Jill is 3 times as old as her cousin Dennis. In 4 years, she will be twice as old as Dennis. What will the sum of their ages be in 4 years?
Answer by stanbon(75887)

Jill is 3 times as old as her cousin Dennis. In 4 years, she will be twice as old as Dennis.
Equations:
j = 3d
j+4 = 2(d+4)
------
Substitute for "j" and solve for "d"::
3d+4 = 2d+8
d = 4 (Dennis's age now)
---
j = 3d = 12 years (Jill's age now)
------------------------------------------
What will the sum of their ages be in 4 years?
Ans: (4+4)+(12+4) = 24
===========================
Cheers,
Stan H.